OpenStudy (anonymous):

If the cost, C, for manufacturing x units of a certain product is given by C=x^2+10+40 find the number of units manufactured at a cost of $7240.

OpenStudy (anonymous):

7240 = x^2 + 10 + 40 simplify 7240 = x^2 + 50 simplify 7190 = x^2 simplify x = 84.8, but since we cant have 0.8 of a product, they only produced 84

OpenStudy (anonymous):

84 was no the right answered.... it was 80

OpenStudy (anonymous):

are you sure you got the original equation right then?

OpenStudy (anonymous):

I have another problem like that one...... to work......... I've be on this type of problem for over an hour now and can't not get the right answer

OpenStudy (anonymous):

it's weird that the equation you gave had 10+40, i suspect there may be an x in there somewhere, most equations dont leave 2 constants on the same side unsimplified like that

OpenStudy (anonymous):

here's the next one.... if the cost, C, for manufacturing x units of a certain product is given by C=x^2-5x+70, find the number of units manufactured at a cost of $8620

OpenStudy (anonymous):

8620 = x^2-5x+70 simplify x^2-5x-8550 = 0 quadratic formula to solve x = 95